OpenStudy (anonymous):

what is the acceleration of a car moving along a straight line path that increases its speed from zero to 100 km/h in 10s ?

OpenStudy (anonymous):

first convert the speed to m/s 100km/h = 100000m/h = 1666.67 m/min. = 27.8 m/s a=(v1-v0)/t a=(27.8m/s - 0 m/s)/10s = 2.78 m/s/s